OpenStudy (anonymous):

David is rowing a boat upstream. The river is flowing at a speed of 2 miles per hour. David starts rowing at a speed of 6 miles per hour, but as he gets tired his speed decreases (at a rate of 1 mile per hour, every hour). Which equation represents the speed of the boat for x hours spent rowing?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@dmezzullo

OpenStudy (dmezzullo):

Do you have any options?

OpenStudy (anonymous):

y = x + 4 y = x − 4 y = 4 − x y = 6 − x y = 6 − 2x

OpenStudy (dmezzullo):

I wanna say the 2nd to last one, Because he's moving at 6 miles then he starts decreasing but doesn't say how long he decreses for.

OpenStudy (anonymous):

thank you

OpenStudy (dmezzullo):

I'm not saying i'm totally right